OpenStudy (lgbasallote):

i was given \[\mathcal L \lbrace 3e^{4t} - e^{-2t} \rbrace\] i got the right answer \[\frac{2s + 10}{(s-4)(s+2)}\] but it lacked the limits that was supposed to be s>4 so my question is why s>4? how do i find that?

OpenStudy (turingtest):

well, you see why not s=4 hopefully any by definition of the laplace transform s>0 so the question is why is 4>s>0 also not admitted, right?

OpenStudy (lgbasallote):

wait...the laplace transform is ALWAYS s>0?

OpenStudy (turingtest):

yes\[\mathcal L\{f(x)\}=\int_0^\infty e^{-st}f(x)dt\]provided that \(s>0\) if somehow the exponent on the integral is not negative it will not converge therefor whatever the exponent is, the bounds are such that it must remain negative for the general case with an unknown function the condition \(s>0) is enough to ensure that the integral converges, but....

OpenStudy (lgbasallote):

ohhh so why isnt it 0<s<4

OpenStudy (turingtest):

you have a specific function of the form \(f(x)=e^{at}\) so our transform becomes\[\int_0^\infty e^{at}\cdot e^{-st}dt=\int_0^\infty e^{(a-s)t}dt\]now again, this integral will only converge if the exponent is negative, hence we have that\[a-s<0\implies s>a\]so you have to look to the integral to find the bounds

OpenStudy (lgbasallote):

so how do i figure out what a is?

OpenStudy (turingtest):

plug in your function when it is of the form e^a

OpenStudy (experimentx):

should be anything ... greater than zero. else your integral is not going to converge.

OpenStudy (lgbasallote):

what do you mean plug in your function when it is of the form e^a? o.O and @experimentX how would i know htat it's 4 and not -2?

OpenStudy (lgbasallote):

i mean how would i know that it's 4?

OpenStudy (turingtest):

whenever you have an exponential function it is easy to check you have two separate ones on yours:\[\int_0^\infty(3e^{4t})(e^{-st})dt+\int_0^\infty(e^{-2t})(e^{-st})dt\]\[3\int_0^\infty e^{(4-s)t}dt+\int_0^\infty e^{-(2+s)t}dt\]so far do you agree?

OpenStudy (lgbasallote):

hmm yeah...

OpenStudy (experimentx):

you can view integration just as sums!! if what would be sum up to infinity if 4-s is negative what would be sum if 4-s is positive?

OpenStudy (turingtest):

we cannot allow either integral to diverge. we achieve that my making sure the exponents in each integral are always negative

OpenStudy (lgbasallote):

so it becomes s > 4 and s>-2 but since it has to be s> 0 it's just s4?

OpenStudy (lgbasallote):

i mean s>4

OpenStudy (experimentx):

... s is assumed to be positive.

OpenStudy (turingtest):

just look at the second integral's exponent: \(-(2+s)t\) since \(s>0\) we know that \(-(2+s)<0\) hence that integral will always converge we need only worry about the other...

OpenStudy (lgbasallote):

hmm i think im beginning to get the idea

OpenStudy (turingtest):

so we say "ok we want the other integral to converge too, so we need the exponent on that guy to be negative" that means 4-s<0 s>4

OpenStudy (experimentx):

i still lack insight on Laplace transform.

OpenStudy (turingtest):

me too, I only have a pdf's worth of understanding though it makes solving some DE's a breeze!

OpenStudy (experimentx):

from analyst point of view ... these things are quite complicated. but somewhere i remember reading Laplace transforms gives you the frequency view. I;ve also seen Laplace transforms used to solve definite integral apart from solving IVP DE

OpenStudy (experimentx):

\[ \Large \int_0^\infty {\sin x \over x}dx\]